#1. Part a. Which of the following bonds is different than the other? (choose one) (-45,...

#1.

Part a. Which of the following bonds is different than the other? (choose one)

  1. (-45, 7.5, 7.5, 7.5, 50)
  2. (-18, 3, 3, 3, 21)
  3. (-6, 1, 1, 1, 7)
  4. (-90, 15, 15, 15, 105)

Part b. Which of the folliwng cash flow streams represent bonds that are trading at par? (check all that apply)

  1. (-100, 5, 5, 105)
  2. (-10, 1, 1, 12)
  3. (-20, 1, 1, 21)
  4. (-50, 2, 2, 50)

Part c. Which of the following must be true regarding the bond described by the cash flow stream (-100, 5, 105)? (Check all that apply)

  1. The yield on the bond is 5%
  2. The bond is trading at par
  3. The coupon rate is 5%
  4. It is equivalent to the bod (-50, 3, 53)
Answer #1

Answer # 1)

a)

The correct option is first ( - 45, 7.5,7.7,7.5, 50) . This is different from the remaining three as the middle three values are 7.5 which is not equal to 50 - 45 = 5

In other cases

Option 2) 21 - 18 = 3
Option 3 ) 7 -6 = 1
Option 4) 105 - 95 = 15.

b)

The correct option is first option

We can see that the value 105 - 100 = 5

The par values are usuually $ 100 or $ 1000 . som municipal bonds trade at $ 5000.

The other options do not have these values.

c)

The correct options are second and third options.

The bond is trading at Par and the coupon rate is 5 %